OpenStudy (anonymous):

Establish cscu-cotu=sinu/(1+cosu)

OpenStudy (anonymous):

if you replace \(\cos(x)\) by \(a\) and \(\sin(x)\) by \(b\) then the left hand side is \[\frac{1}{b}-\frac{a}{b}\]

OpenStudy (anonymous):

subtracting gives you \[\frac{1-a}{b}\] or \[\frac{1-\cos(x)}{\sin(x)}\]

OpenStudy (anonymous):

so I get \[\frac{ cosu - \sin ^2u }{ cosusinu } = \frac{ sinu }{ 1+cosu }\]

OpenStudy (anonymous):

now lets see if we can show \[\frac{1-\cos(x)}{\sin(x)}=\frac{\sin(x)}{1+\cos(x)}\]

OpenStudy (anonymous):

how did you get the left side?

OpenStudy (anonymous):

in your subtraction you overcomplicated it you can use just \(\sin(x)\) as our denominator

OpenStudy (anonymous):

\[\frac{1}{\sin(x)}-\frac{\cos(x)}{\sin(x)}=\frac{1-\cos(x)}{\sin(x)}\]

OpenStudy (anonymous):

oh I got it thanks!

OpenStudy (anonymous):

now how to get \[\frac{1-\cos(x)}{\sin(x)}=\frac{\sin(x)}{1+\cos(x)}\]? one hint is that \[\sin^2(x)+\cos^2(x)=1\] or \[\sin^2(x)=1-\cos^2(x)\] factoring on the left gives \[\sin^2(x)=(1+\cos(x))(1-\cos(x))\]

OpenStudy (anonymous):

or if you like \[\sin(x)\times \sin(x)=(1+\cos(x))(1-\cos(x))\] divide both sides by \(1+\cos(x)\) and divide by \(\sin(x)\) and you get it \[\frac{\sin(x)}{1+\cos(x)}=\frac{1-\cos(x)}{\sin(x)}\]

OpenStudy (anonymous):

i meant "factoring on the right"

OpenStudy (anonymous):

so the answer is \[\frac{ 1-\cos^2u }{ sinu(1+cosu) }\]

OpenStudy (anonymous):

for the left side

OpenStudy (anonymous):

hmm no

OpenStudy (anonymous):

you good to this line \[\sin(x)\times \sin(x)=(1+\cos(x))(1-\cos(x))\]

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

ok i think i have confused you i am not starting with \[\frac{\sin(x)}{1+\cos(x)}=\frac{1-\cos(x)}{\sin(x)}\] but i am going to end with it

OpenStudy (anonymous):

but if you want to start with that, you can "cross multiply" and see that it is true that is not really a proof, but it might do

OpenStudy (anonymous):

\[\sin^2(x)+\cos^2(x)=1\\ \sin^2(x)=1-\cos^2(x)\\ \sin(x)\times \sin(x)=(1+\cos(x))(1-\cos(x))\] you can get from that line to what you want to prove, by dividing both sides by \(\sin(x)\) and by \(1+\cos(x))\)

OpenStudy (anonymous):

thanks !

OpenStudy (anonymous):

yw hope i did not confuse you too much